# Large Language Model Unlearning for Source Code

### v1

@@ V-C5 RQ5: Perceptual Quality Evaluation
Following the work [48, 49], we conduct a perceptual quality evaluation to assess the subjective human experience of our approach.
@@
Given that human evaluation can only assess a limited portion of samples due to practical constraints, and recent research suggests that LLMs can substitute human evaluation to some extent with high evaluation consistency [50], we introduced GPT-4 to evaluate all samples across all downstream tasks.
TABLE II: Win rate comparison on perceptual evaluation.
@@
Results. The results of the perceptual quality evaluation are shown in Figure II. We find that our approach achieves win rates above 70% compared to baselines. According to the reasons provided by human evaluators, the common feedback was that our approach still produced valuable outputs after forgetting specific content, while baselines often produced empty outputs or messy code.
VI Threats to Validity
We have taken all reasonable steps to mitigate potential threats that could hamper the validity of this work. Specifically, we identify and attempt to mitigate two main categories of threats: internal validity threats and external validity threats. Internal threats to validity pertain to factors that may impact the observed causal relationships in our experiments. A potential threat arises from the sensitivity of LLM unlearning approaches to hyperparameter settings, especially learning rates. To mitigate this, we conduct extensive hyperparameter tuning through grid search for learning rates and report by averaging multiple results. We kept other hyperparameters consistent with established practices from prior studies to minimize unintended biases. External threats to validity concern the generalizability of our findings beyond the experimental contexts considered in this study. We address this by evaluating PROD on three representative unlearning tasks, each reflecting distinct practical scenarios encountered in real-world software engineering contexts. Additionally, we validate our approach across multiple LLM architectures, including CodeLlama, Deepseek-Coder, Qwen2.5-Coder, and Starcoder, which vary in their training data, model architectures, and parameter scales. Thus, we believe the concern about the generalizability of our findings is mitigated.
VII Conclusion
In this paper, we have investigated existing LLM unlearning approaches on source code, identifying significant utility degradation. To this end, we propose PROD, a novel unlearning approach that reconstructs the output distribution by suppressing target code and preserving essential code-generation capabilities. We also establish a benchmark covering copyrighted code, insecure code, and deprecated API unlearning tasks. Extensive experiments show PROD significantly outperforms existing methods in both forgetting performance and user experience, while also exhibiting broad applicability and robustness against unlearning attacks.

### v2

@@ Appendix C Perceptual Quality Evaluation
Following the work (macketanz-etal-2022-perceptual; DBLP:conf/icmcs/ZhangLSLMZ23), we conduct a perceptual quality evaluation to assess the subjective human experience of our approach.
@@
Given that human evaluation can only assess a limited portion of samples due to practical constraints, and recent research suggests that LLMs can substitute human evaluation to some extent with high evaluation consistency (liu-etal-2023-g), we introduced GPT-4 to evaluate all samples across all downstream tasks.
Table 2: Win rate comparison on perceptual evaluation.
@@
Results. The results of the perceptual quality evaluation are shown in Figure 2. We find that our approach achieves win rates above 70% compared to baselines. According to the reasons provided by human evaluators, the common feedback was that our approach still produced valuable outputs after forgetting specific content, while baselines often produced empty outputs or messy code.
Appendix D Extended Related Work
Code Generation with LLMs
Code generation is the computational process of automatically producing executable source code from specifications or natural language inputs. Generative models such as ChatGPT (achiam2023gpt4) and Claude (claude3.5) have demonstrated strong capabilities in code generation, benefiting from pretraining on large-scale data and code repositories. Given the immense potential of LLMs, researchers have developed code-specialized LLMs, such as CodeX (codex), CodeLlama (CodeLlama), DeepSeek-Coder (zhu2024deepseekcoder), Qwen2.5 Coder (hui2024qwen2.5-coder), and CodeGemma (team2024codegemma), which further enhanced LLMs’ capabilities across various software development tasks. However, several critical challenges remain unresolved in leveraging LLMs for code generation, including concerns about safety, legality, and reliability.
Recently, an emerging concern is that LLM-generated code may inadvertently violate security best practices. Because LLMs learn from massive public code repositories, many of which include buggy or outdated patterns, they can propagate known vulnerabilities or poor coding habits. Researchers have begun exploring security-focused code generation techniques (safecoder; mohsin2024can; wu2024versicode; wang2024and; xu2024first; fu2023security). Representative approaches include SaferCode (safecoder), which augments LLM training with carefully curated secure code datasets, and work (mohsin2024can), which proposes incorporating security examples directly into prompts to avoid unsafe coding patterns. Beyond security issues, VersiCode (wu2024versicode) introduces a version-specific benchmark to evaluate how models handle API changes across different releases, revealing that version-controllable code generation poses significant challenges. Failing to adapt can lead to functional errors and long-term maintenance issues. Similarly, SecuCoGen (wang2024and) systematically analyzes deprecated API usage in LLM-based code completion. Testing seven advanced models on over 28k prompts and 145 API changes, they observe a substantial rate of outdated suggestions, indicating that many LLMs fail to migrate to newer APIs. Additionally, copyright concerns have also been raised regarding LLM-generated code. LiCoEval (xu2024first) systematically evaluates LLMs’ license compliance and finds that models sometimes produce code fragments that closely resemble copyrighted implementations, often without proper attribution.
LLM Unlearning
Machine unlearning, the process of removing specific data and its influence from a trained machine learning model, has garnered significant attention in recent years. Since (DBLP:conf/sp/CaoY15) first introduced the concept of machine unlearning, laying the groundwork for subsequent research. Due to the widespread application and the relative simplicity of classification models, they become the focal point for most of the machine unlearning works (DBLP:conf/cvpr/GolatkarAS20; DBLP:journals/corr/abs-2002-02730; DBLP:conf/aistats/IzzoSCZ21; DBLP:conf/sigmod/SchelterGD21; DBLP:conf/sp/BourtouleCCJTZL21; DBLP:conf/nips/SekhariAKS21). Machine unlearning approaches can be broadly categorized into three types: data-reversed training (DBLP:journals/tnn/TarunCMK24; DBLP:conf/infocom/LiuFCLMWM22; DBLP:journals/tifs/ChundawatTMK23), influence function based approaches (DBLP:conf/aistats/IzzoSCZ21), and optimization based unlearning (DBLP:conf/icml/GuoGHM20; DBLP:conf/alt/Neel0S21). Data-reversed training techniques employ noise injection to reverse or weaken the model’s memory and dependence on specific data. Influence function-based approaches evaluate the impact of training samples on the model by perturbing model parameters, which are less common in LLM unlearning due to the computational complexity of Hessian matrix inversion. Optimization-based unlearning, which is the focus of our work, achieves forgetting through direct model optimization. The rapid advancement of LLMs presents more significant challenges for machine unlearning. Their massive parameter scales and generative nature make traditional unlearning approaches unsuitable (DBLP:journals/natmi/LiuYJCBHYLXLVBKL25). Recent research efforts have been dedicated to the domain of LLM unlearning within the field of NLP, proposing methods such as GA (Unlearning), DPO (DPO), NPO (NPO), and FLAT (FLAT), which aim to address forgetting tasks in the field of NLP, such as user privacy unlearning, copyrighted books unlearning, and harmful content unlearning (Unlearning). To facilitate the evaluation of LLM unlearning, several benchmarks have been proposed. These include synthetic profiles and question-answer pairs about fictitious authors (DBLP:journals/corr/abs-2401-06121), as well as the forgetting of specific content like Harry Potter books (DBLP:journals/corr/abs-2310-10683). Moreover, to assess the vulnerabilities of unlearned models, various attack approaches have been employed, such as prefix injections (Jailbroken; DBLP:journals/corr/abs-2408-10682; qi2025safety). These attacks help researchers understand the robustness of unlearning techniques and identify areas for improvement.